The number of [H+] ions in a solution is 3.3. Calculate the number of [OH-] ions. _ (use E to signify x10)
1
Expert's answer
2019-05-20T08:24:44-0400

If 3.3 Is -log[H+]

Then -log[OH-] =14 -3.3 = 10.7

[OH-] =10E-10.7


2. If 3.3 Is n(H+)

Then

C(H+) = 3.3 mol/dm3

-log[OH-]= 14 + log(C(H+))

-log[OH-]= 14.519

[OH-]= 3.0303E-15